Todd is able to pay $360 a month for 6 years to finance a car purchase.

a. If the interest rate is 6.7 percent compounded monthly, how much can Todd afford to borrow to buy a car?

b. What is the effective annual rate of Todd's loan?
Answer #1

Per month payment =360
Number of months =6*12 =72
Rate per month =6.7%.12
a. Amount Todd can afford to borrow =PMT*(((1-(1+r)^-n)/r) =360*(((1-(1+6.7%/12)^-72)/(6.7%/12)) =21295.08
b. Effective rate=(1+APR/12)^12-1 =(1+6.7%/12)^12-1 =6.91%